# Neyman-Pearson multiclass classification under label noise via empirical likelihood

### 23 Mar 2026 v1

@@ 11 $(1,g^{\top}(X_{n}))^{\top}$ $\omega_{n1}^{(t)}$
Although effective for controlling type I error, Yao et al. (2023) showed that the NP umbrella algorithm becomes too conservative when the training data are contaminated, i.e., when labels are corrupted as in Assumption 2.1. This is because ignoring label noise in the NP umbrella algorithm only guarantees $\widetilde{P}_{0}^{*}(\phi\neq 0)\leq\alpha$ instead of $P_{0}^{*}(\phi\neq 0)\leq\alpha$. To address this issue, they extended the NP umbrella algorithm to handle the noisy-label setting as follows:
@@
This approach requires knowledge of the parameters $m_{0}^{*}={\mathbb{P}}(Y=0|\widetilde{Y}=0)$ and $m_{1}^{*}={\mathbb{P}}(Y=0|\widetilde{Y}=1)$ in the transition matrix; we refer to this method as NPC. When these quantities are unknown, Yao et al. (2023) further extended the method to the case where only a lower bound on $m_{1}^{*}$ and an upper bound on $m_{0}^{*}$ are available. In this approach, they replace $m_{0}^{*}$ and $m_{1}^{*}$ with their bounds in the last step. We refer to this version as NPC.
@@
[…]
The Dry Bean dataset contains $13,611$ instances of dry beans along with their corresponding types (Koklu and Ozkan 2020). Each instance is described by $16$ features capturing the beans’ physical properties, derived from shape, size, and color measurements. Specifically, the features include $11$ geometric, $1$ shape-related, and $2$ color-related attributes. The dataset contains seven types of dry beans with the following sample sizes: Barbunya ($1,322$), Bombay ($522$), Cali ($1,630$), Dermosan ($3,546$), Horoz ($1,928$), Seker ($2,027$), and Sira ($2,636$). For convenience, we recode the types into classes 0 through $6$. The prediction task is to classify the type of dry bean based on its $16$ physical features. The dataset is publicly available from the UCI Machine Learning Repository at this link.
@@
We consider the following NPMC problem from Tian and Feng (2025):
$\begin{split}\min_{\phi}&\quad\frac{1}{4}\left[P_{2}^{*}(\{\phi(X)\neq 2\})+P_{4}^{*}(\{\phi(X)\neq 4\})+P_{5}^{*}(\{\phi(X)\neq 5\})+P_{6}^{*}(\{\phi(X)\neq 6\})\right]\\ \text{s.t.}&\quad P_{0}^{*}(\{\phi(X)\neq 0\})\leq 0.05,\quad P_{1}^{*}(\{\phi(X)\neq 1\})\leq 0.01,\quad P_{3}^{*}(\{\phi(X)\neq 3\})\leq 0.03.\end{split}$
@@
[…]
@@
[…]
@@
${\mathbb{E}}\left\{\frac{\partial h(\mbox{$\nu$}^{*},\mbox{$\theta$}^{*})}{\partial\nu_{j}}\right\}=0\quad\text{and}\quad{\mathbb{E}}\left\{\frac{\partial h(\mbox{$\nu$}^{*},\mbox{$\theta$}^{*})}{\partial w_{j}}\right\}=0$
[…]
@@
[…]
@@
${\bm{v}}=(\mbox{$\nu$}^{\top},\mbox{$\theta$}^{\top})^{\top}\quad\text{and}\quad{\bm{v}}^{*}=((\mbox{$\nu$}^{*})^{\top},(\mbox{$\theta$}^{*})^{\top})^{\top}\quad\text{and}\quad\widehat{\bm{v}}=(\widehat{}\mbox{$\nu$}^{\top},\widehat{}\mbox{$\theta$}^{\top})^{\top}.$
@@
Following the proofs of Lemma 1 and Theorem 1 of Qin and Lawless (1994) and the proof of Theorem 2 in Qin et al. (2015), we obtain
@@
[…]
@@
$\frac{d\widetilde{P}_{l}^{*}}{dP_{0}^{*}}(x)=\sum_{k=1}^{K}M_{lk}^{*}\exp\big(\gamma_{k}^{*}+\langle\beta_{k}^{*},g(x)\rangle\big),\quad\forall l\in[K],$ (48)
@@
[…]
@@
[…]
@@
$F(tv)=\sum_{k=1}^{K}c_{k}\exp(t\langle\beta_{k},v\rangle)=\sum_{k=1}^{K}c_{k}\exp(\lambda_{k}t)=0,\quad\forall t\in\mathbb{R}.$
@@
$\sum_{k=1}^{K}M_{lk}f_{k}(x)p_{0}(x)=\sum_{j=1}^{K}\widetilde{M}_{lj}\widetilde{f}_{j}(x)\widetilde{p}_{0}(x),\quad\forall l\in[K].$
@@
[…]
@@
$\qquad-\sum_{m=1}^{K}\Big[A_{km}\exp(\widetilde{\gamma}_{m})\Big]\exp\big(\langle\widetilde{\beta}_{m},g(x)\rangle\big)=0.$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
See Wainwright (2019, Example 5.24). ∎
[…]
@@
[…]
See Wainwright (2019, Theorem 4.10). ∎
[…]
@@
$\|h-\tilde{h}\|_{n}\leq\delta,\quad\|k-\tilde{k}\|_{n}\leq\delta.$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Using standard bounded ball covering number results (see Wainwright (2019, Lemma 5.7) for example), the $\epsilon$-covering of a radius $B$ ball in ${\mathbb{R}}^{d}$ is at most
@@
By Vershynin (2018, Theorem 8.3.18), there exists some universal constant $C>0$ such that
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\widehat{}\mbox{$\lambda$}\overset{a.s.}{\longrightarrow}\mbox{$\lambda$}^{*}\quad\text{as}\quad n\to\infty.$
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Using standard bounded ball covering number results (see Wainwright (2019, Lemma 5.7) for example), the $\epsilon$-covering of a radius $B$ ball in ${\mathbb{R}}^{(K-1)(d+3)}$ is at most
@@
[…]
@@
[…]

### 21 Apr 2026 v2

@@ B.2.3 NP umbrella type control
Although effective for controlling type I error, Yao et al., (2023) showed that the NP umbrella algorithm becomes too conservative when the training data are contaminated, i.e., when labels are corrupted as in Assumption 2.1. This is because ignoring label noise in the NP umbrella algorithm only guarantees $\widetilde{P}_{0}^{*}(\phi\neq 0)\leq\alpha$ instead of $P_{0}^{*}(\phi\neq 0)\leq\alpha$. To address this issue, they extended the NP umbrella algorithm to handle the noisy-label setting as follows:
@@
This approach requires knowledge of the parameters $m_{0}^{*}={\mathbb{P}}(Y=0|\widetilde{Y}=0)$ and $m_{1}^{*}={\mathbb{P}}(Y=0|\widetilde{Y}=1)$ in the transition matrix; we refer to this method as NPC. When these quantities are unknown, Yao et al., (2023) further extended the method to the case where only a lower bound on $m_{1}^{*}$ and an upper bound on $m_{0}^{*}$ are available. In this approach, they replace $m_{0}^{*}$ and $m_{1}^{*}$ with their bounds in the last step. We refer to this version as NPC.
@@
B.2.4 Real data experiment: dry bean dataset
The Dry Bean dataset contains $13,611$ instances of dry beans along with their corresponding types (Koklu and Ozkan, , 2020). Each instance is described by $16$ features capturing the beans’ physical properties, derived from shape, size, and color measurements. Specifically, the features include $11$ geometric, $1$ shape-related, and $2$ color-related attributes. The dataset contains seven types of dry beans with the following sample sizes: Barbunya ($1,322$), Bombay ($522$), Cali ($1,630$), Dermosan ($3,546$), Horoz ($1,928$), Seker ($2,027$), and Sira ($2,636$). For convenience, we recode the types into classes 0 through $6$. The prediction task is to classify the type of dry bean based on its $16$ physical features. The dataset is publicly available from the UCI Machine Learning Repository at this link.
@@ B.2.4 Real data experiment: dry bean dataset
We consider the following NPMC problem from Tian and Feng, (2025):
$\begin{split}\min_{\phi}&\qquad\frac{1}{4}\left[P_{2}^{*}(\{\phi(X)\neq 2\})+P_{4}^{*}(\{\phi(X)\neq 4\})+P_{5}^{*}(\{\phi(X)\neq 5\})+P_{6}^{*}(\{\phi(X)\neq 6\})\right]\\ \text{s.t.}&\qquad P_{0}^{*}(\{\phi(X)\neq 0\})\leq 0.05,\qquad P_{1}^{*}(\{\phi(X)\neq 1\})\leq 0.01,\qquad P_{3}^{*}(\{\phi(X)\neq 3\})\leq 0.03.\end{split}$
@@ C.2 Derivation of the profile log empirical likelihood
C.3 Asymptotic normality of EL based parameter estimation
C.3.1 Notation and preparation
@@ C.3.2 Expectation of score function at truth
We now proceed with the proof of Lemma C.1.
Proof of Lemma C.1.
@@
${\mathbb{E}}\left\{\frac{\partial h(\mbox{$\nu$}^{*},\mbox{$\theta$}^{*})}{\partial\nu_{j}}\right\}=0\qquad\text{and}\qquad{\mathbb{E}}\left\{\frac{\partial h(\mbox{$\nu$}^{*},\mbox{$\theta$}^{*})}{\partial w_{j}}\right\}=0$
for all $j$. The remaining equalities follow similarly from Lemma C.2.
@@
where the last equality follows from Lemma C.2. From the exponential tilting relationship in (6), we have:
@@ C.3.3 Asymptotic normality of the estimator
${\bm{v}}=(\mbox{$\nu$}^{\top},\mbox{$\theta$}^{\top})^{\top}\qquad\text{and}\qquad{\bm{v}}^{*}=((\mbox{$\nu$}^{*})^{\top},(\mbox{$\theta$}^{*})^{\top})^{\top}\qquad\text{and}\qquad\widehat{\bm{v}}=(\widehat{}\mbox{$\nu$}^{\top},\widehat{}\mbox{$\theta$}^{\top})^{\top}.$
@@
Following the proofs of Lemma 1 and Theorem 1 of Qin and Lawless, (1994) and the proof of Theorem 2 in Qin et al., (2015), we obtain
@@
Using the central limit theorem, Lemma C.1, and Slutsky’s theorem, we have
@@ C.4 Identifiability
$\frac{d\widetilde{P}_{l}^{*}}{dP_{0}^{*}}(x)=\sum_{k=1}^{K}M_{lk}^{*}\exp\big(\gamma_{k}^{*}+\langle\beta_{k}^{*},g(x)\rangle\big),\qquad\forall l\in[K],$ (46)
@@ C.4.1 Assumptions
C.4.2 Supporting Lemma
Assumption C.2 is crucial to ensure that the exponential functions in our mixture do not collapse into linear dependence over the data support.
Lemma C.3.
Under Assumption C.2, for any finite set of pairwise distinct vectors $\beta_{1},\dots,\beta_{K}\in\mathbb{R}^{d}$, the functions $f_{k}(x)=\exp(\langle\beta_{k},g(x)\rangle)$ are linearly independent over $\mathcal{X}$}. That is, if
@@ C.4.2 Supporting Lemma
Let $z=g(x)$. By hypothesis, $F(z)=\sum_{k=1}^{K}c_{k}\exp(\langle\beta_{k},z\rangle)=0$ holds for all $z\in{\mathcal{Z}}$, where ${\mathcal{Z}}=\{g(x):x\in{\mathcal{X}}\}\subseteq\mathbb{R}^{d}$ is the image of $\mathcal{X}$}. Because $\mathcal{Z}$} contains a non-empty open set (Assumption C.2) and $F(z)$ is real-analytic on $\mathbb{R}^{d}$, the identity theorem for real-analytic functions dictates that $F(z)=0$ for all $z\in\mathbb{R}^{d}$.
@@
$F(tv)=\sum_{k=1}^{K}c_{k}\exp(t\langle\beta_{k},v\rangle)=\sum_{k=1}^{K}c_{k}\exp(\lambda_{k}t)=0,\qquad\forall t\in\mathbb{R}.$
@@ C.4.3 Proof of identifiability
$\sum_{k=1}^{K}M_{lk}f_{k}(x)p_{0}(x)=\sum_{j=1}^{K}\widetilde{M}_{lj}\widetilde{f}_{j}(x)\widetilde{p}_{0}(x),\qquad\forall l\in[K].$
@@
By Assumption C.3, $\bm{M}$} is invertible. Left-multiplying by ${\bm{M}}^{-1}$ and defining $A={\bm{M}}^{-1}\widetilde{{\bm{M}}}$, we obtain $f(x)=A\widetilde{f}(x)R(x)$. The $k$-th component is:
@@
$\hskip 18.49988pt-\sum_{m=1}^{K}\Big[A_{km}\exp(\widetilde{\gamma}_{m})\Big]\exp\big(\langle\widetilde{\beta}_{m},g(x)\rangle\big)=0.$
@@
where $W_{v}$ is the net scalar coefficient for the unique vector $v$. Because the elements of $V$ are now pairwise distinct by definition, Lemma C.3 directly applies, dictating that $W_{v}=0$ for all $v\in V$.
Suppose there is an “active” index $j$ on the left side where $A_{1j}\neq 0$. The exponent vector for this term is $v^{*}=\beta_{k}+\widetilde{\beta}_{j}$. Because the original vectors $\{\widetilde{\beta}_{j}\}_{j=1}^{K}$ are pairwise distinct (Assumption C.1), the shifted vectors $\{\beta_{k}+\widetilde{\beta}_{j}\}_{j=1}^{K}$ are also strictly pairwise distinct. This means no other term in the first sum shares the exponent vector $v^{*}$.
@@
Suppose the permutation is not the identity, meaning there is some $k$ where $\sigma^{-1}(k)\neq k$. Then ${\bm{M}}^{-1}]_{\sigma^{-1}(k),k}$ is an off-diagonal entry. By Assumption C.3, all off-diagonal entries of ${\bm{M}}^{-1}$ are nonpositive ($\leq 0$). This implies that $(\widetilde{{\bm{M}}})^{-1}]_{k,k}\leq 0$, which explicitly contradicts the requirement in Assumption C.3 that $(\widetilde{{\bm{M}}})^{-1}$ must have strictly positive diagonal entries.
@@ C.5.1 Definitions
C.5.2 Lemmas
Lemma C.4 (Bounds for Vapnik–Chervonenkis classes).
@@ C.5.2 Lemmas
Proof of Lemma C.4.
See Wainwright, (2019, Example 5.24). ∎
Lemma C.5 (Uniform law via Rademacher complexity).
@@
where $\mathfrak{R}_{n}({\mathcal{F}})$ is the Rademacher complexity of $\mathcal{F}$} defined in Definition C.1.
Proof of Lemma C.5.
See Wainwright, (2019, Theorem 4.10). ∎
Lemma C.6 (Covering number of product of uniformly bounded functions).
@@
$\|h-\tilde{h}\|_{n}\leq\delta,\qquad\|k-\tilde{k}\|_{n}\leq\delta.$
@@
Lemma C.7 (Covering number of pointwise maximum functions).
@@
C.6 Guaranteed type I error control under binary classification
@@
C.6.1 Proof of type I error under binary classification
@@
Since $\mathcal{F}$} is uniformly bounded by 1, by Lemma C.5, we have with probability at least $1-\delta$:
@@
where $\mathfrak{R}_{n}({\mathcal{F}})$ is the Rademacher complexity of $\mathcal{F}$}. We have shown in Lemma C.9 that
@@
where $C^{\prime}$ is the universal constant specified in Lemma C.9. Combining all bounds, we conclude that:
@@
C.6.2 Rademacher complexity of weighted logistic classifiers
Lemma C.8 (Covering number of weighted logistic threshold functions).
Let $\|\cdot\|_{n}$ denote the empirical $L_{2}$-norm defined in Lemma C.6. Consider the function class
@@ C.6.2 Rademacher complexity of weighted logistic classifiers
Since both $\mathcal{H}$} and $\mathcal{K}$} are uniformly bounded by $1$, Lemma C.6 implies
@@
Using standard bounded ball covering number results (see Wainwright, (2019, Lemma 5.7) for example), the $\epsilon$-covering of a radius $B$ ball in ${\mathbb{R}}^{d}$ is at most
@@
By Vershynin, (2018, Theorem 8.3.18), there exists some universal constant $C>0$ such that
@@
Lemma C.9 (Rademacher complexity of weighted binary classifier).
Let $\mathcal{F}$} be the same function class defined in Lemma C.8. Assume that $\mathbb{E}\bigl[\|g(X)\|_{2}^{2}\bigr]\leq R^{2}$ for some $R>0$, we have
@@
where $C^{\prime}=24\sqrt{\pi}(1+2BR)C$ and $C$ is the universal constant in Lemma C.8.
Proof of Lemma C.9.
Note that we have $\sup_{f,g\in{\mathcal{F}}}\|f-g\|_{n}=1$. By Lemma C.4, we have
@@
C.7 Guaranteed class specific error control under multiclass classification
C.7.1 Proof of class-specific error control
@@
Almost surely, there are no ties among $\{Z_{k}\}_{k=0}^{K-1}$, that is, $\eta>0$ where $\eta$ is the gap between the largest and second largest entries in this set. By Lemma C.11 and Theorem 3.2, we have
@@
C.7.2 Asymptotic convergence of dual problem
Lemma C.10 (Uniform convergence of dual objective).
@@ C.7.2 Asymptotic convergence of dual problem
Bound $T_{1}$: Since $\mathcal{F}$} is uniformly bounded and satisfies the conditions of Lemma C.5, we have with probability at least $1-\delta$:
@@
where $\mathfrak{R}_{n}({\mathcal{F}})$ is the Rademacher complexity of $\mathcal{F}$}. By Lemma C.13, there exists a constant $C^{\prime}>0$ such that:
@@
$\widehat{}\mbox{$\lambda$}\overset{a.s.}{\longrightarrow}\mbox{$\lambda$}^{*}\qquad\text{as}\qquad n\to\infty.$
Proof of Lemma C.11.
@@
where the strict inequality follows from $\nabla^{2}G(\mbox{$\lambda$}^{*})\prec 0$, which guarantees $\mbox{$\lambda$}^{*}$ is the unique maximizer of $G$ in $\bar{B}_{\epsilon}(\mbox{$\lambda$}^{*})$. By Lemma C.10, $\sup_{\mbox{$\lambda$}\in\overline{B}_{\epsilon}(\mbox{$\lambda$}^{*})}|\widehat{G}(\mbox{$\lambda$})-G(\mbox{$\lambda$})|\to 0$ almost surely as $n\to\infty$. Thus, (50) holds almost surely, completing the proof. ∎
C.7.3 Rademacher complexity of weighted multinomial logistic classifiers
Lemma C.12 (Covering number of weighted multinomial logistic classifier).
Let $\|\cdot\|_{n}$ denote the empirical $L_{2}$-norm defined in Lemma C.6. Let $B>0$ be some fixed constant, $\Lambda\subseteq{\mathbb{R}}_{+}^{|{\mathbb{S}}|}$ be any bounded set, and
@@ C.7.3 Rademacher complexity of weighted multinomial logistic classifiers
Proof of Lemma C.12.
Since $\sup_{f,g\in{\mathcal{F}}}\|f-g\|_{n}$ is bounded by a finite constant, we show the upper bound for the Rademacher complexity of $\mathcal{F}$} via Lemma C.4. As such, we consider the covering number of $\mathcal{F}$}. Denote
@@
where $C_{k}(\lambda,w)=\frac{\rho_{k}+\lambda\mathbbm{1}(k\in{\mathcal{S}})}{w}$. By Lemma C.6, for any $t>0$, we have
@@
Using standard bounded ball covering number results (see Wainwright, (2019, Lemma 5.7) for example), the $\epsilon$-covering of a radius $B$ ball in ${\mathbb{R}}^{(K-1)(d+3)}$ is at most
@@
Lemma C.13 (Rademacher complexity of weighted multinomial logistic classifier).
Let $\mathcal{F}$} be the same as that defined in C.12. Under Assumption 4.2, we have
@@
Proof of Lemma C.13.
Note that we have $\sup_{f,g\in{\mathcal{F}}}\|f-g\|_{n}\leq M$ for some finite constant $M>0$. By Lemma C.4, we have